### v2

# Predictable Drifts in Collective Cultural Attention: Evidence from Nation-Level Library Takeout Data

Abstract
Predicting changes in consumer attention for cultural products, such as books, movies, and songs, is notoriously difficult. Past research on predicting the popularity of individual products suggests the existence of intrinsic prediction limits. However, little is known about the limits for predicting collective attention across cultural products. Here, we analyze four years of nationwide library loan data for approximately 2 million individuals, comprising over 100 million loans of more than 660,000 unique books. We find that culture, as measured by popularity distributions of loaned books, drifts continually from month to month at a near-constant rate, leading to a growing divergence over time, and that drifts vary between different book genres. By linking book loans to registry data, we investigate the influence of age, sex, educational level, and geographical area on cultural drift, finding heterogeneous effects from the different demographic groups. Our findings have important implications for market forecasting and developing robust recommender systems, highlighting the need to account for specific drift dynamics for different types of items and demographic groups.
keywords:
cultural markets, predictability, library data, social systems, data drift
Cultural products drive billion-dollar industries, influence public discourse, and shape our personal and collective identities. Yet, their popularity remains notoriously hard to predict, even despite advances in predictive modeling and the growing availability of large-scale cultural consumption data [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. This unpredictability coincides with extremely unequal distributions of popularity, where a small number of products capture a large share of consumer attention [12, 13, 14, 15, 16]. Seminal work suggests that the unpredictability and inequality observed in cultural markets can be explained by high levels of social influence [17, 18]. Strong social influence on individual decisions can produce information cascades in cultural markets and other social systems, where large numbers of consumers simultaneously attend to the same cultural items [19, 20, 21]. Multiple studies have systematically analyzed the dynamics of collective attention, revealing characteristic patterns in how popular products and major events enter into and fade from collective attention [22, 23, 24, 25, 26, 27].
However, existing approaches have not been able to address important questions about collective consumer attention in cultural markets. How much does collective attention change on aggregate? Are these changes universal across products and consumer groups? Are they predictable? Answers to these questions have implications for our theoretical understanding of cultural markets and for practical applications involving models of market dynamics, for instance, in developing robust recommender systems. Resolving them requires comprehensive cultural consumption data that moves beyond the current focus on hit products as well as an analytical framework for quantifying changes in collective attention at the market level.
Here, we use a unique nation-scale book loan dataset which contains records from all public libraries in Denmark, featuring 100 million book loans from 2 million individuals (over one third of the Danish population), to study how cultural attention evolves at a societal scale. In Denmark, library borrowing has a low barrier to entry since library services are well-funded, free to use, and accessible anywhere in the country. In addition, Danish public libraries have a wide coverage, and any material present in the library system can be ordered from any local library. For these reasons, libraries serve as a primary means of accessing books for Danish citizens. According to the best available data on book sales in Denmark, the number of book loans from public libraries is consistently almost three times the number of book sales (see SI Fig. LABEL:book_sales_loans_comparison based on [28]). Further, Denmark is one of the most digitalized countries in the world, making it possible to link each individual book loan to registry data, which provides information about library users’ demographic characteristics, such as age, sex, education level, zip code, etc. Given the unprecedented scope of this dataset, we can identify changes in collective consumer attention for a wide range of books across a large number of consumers.
Our approach differs fundamentally from prior work on collective attention. Rather than focusing on the individual trajectories of popular items in isolation, we analyze the changes in consumer attention across all items in the market, which we refer to as ‘drifts’ in collective attention. This novel approach to measuring changes in collective attention eliminates the need to filter away less popular products, capturing attention dynamics across the full popularity spectrum. In addition, it requires no assumptions about how individual items gain or lose consumer attention, accommodating diversity in popularity trajectories.
Our analysis is split up into three parts. First, we quantify drift in collective attention, finding that culture has a constant turnover rate across months. Second, we dig into which demographic groups contribute to this drift, finding that age, sex, and place of residence have a large influence. Third, we account for the contributions of each unique book title to collective attention drift. We use our findings to build a simple model that can accurately estimate drift without needing to account for individual product trajectories.
1 Results
We use comprehensive records of all Danish public library loans (excluding university and school libraries) from start 2020 to end 2023. Each entry in our dataset corresponds to the loan of a single book and contains information on the date of the loan, a unique identifier of the book, a pseudonymized loaner ID, the library where the book was loaned, as well as metadata on the books and loaners (see Fig. 1A). The dataset contains more than 100 million unique book loans from 513 libraries, including eBooks from digital services (see SI Fig. LABEL:material_type_distribution for the distribution of material types) and features loans from over 45% of the entire Danish population. A single book title has multiple identifiers for different media (physical book, audiobook, etc.) and editions. Throughout our analysis, we treat each variant of a book as a single entity (see Methods for details on how we identify individual books). We link loaner IDs to national registry data to obtain sociodemographic information, including age, sex, education level, and area of residence for each loaner (see Methods).
Figure 1: Dataset and measurement overview. A, Each entry in the dataset corresponds to the loan of a single book, illustrated as vertical lines. Some loaner attributes (sex and birthdate) are static, whereas others (loaner residence and education level) are dynamic. We aggregate loans into monthly time periods and measure the local (between successive periods) and global (between a given period and a reference period) drifts in cultural attention. B, Aggregate daily number of loans. A nationwide COVID-19 lockdown, effective March 13, 2020, and lasting until May 18, 2020, shut down all libraries. The few loans of physical books during this period can be attributed to some libraries putting out ‘book bags’ for their loaners. A partial lockdown, initiated on December 9, 2020, limited public access to library interiors in 38 municipalities. During these lockdown periods, there was a noticeable decrease in the number of books borrowed from physical libraries and an increase in digital loans. The inset shows an example of the weekly fluctuations in takeout numbers over two weeks from 24/01/2022 to 06/02/2022. C, Weekly number of loans for 20 popular books in the dataset. D, Complementary cumulative distribution function (CCDF) of the number of book loans for the more than 660,000 books over the entire four-year period. The inset shows the CCDFs for three main categories of books. E, Number of loans for different sociodemographic subgroups.
To quantify drifts in library book consumption, we use the Jensen-Shannon Divergence (JSD). The JSD is an information-theoretic measure that is well-suited to measure changes in sequences of frequency distributions [29]. In the social science literature, JSD and related measures have mainly been used to analyze text sequences [30, 31, 32, 33, 34, 35, 36, 37]. In this study, we use it on sequences of monthly book popularity distributions (see Methods for technical details). We use months as the aggregation level since it allows us to capture changes in consumption without being overly sensitive to minor variations in book popularity that do not reflect meaningful changes in consumer attention. If the JSD between loans in two months is 0, it means the popularity distributions are identical; if it is 1, entirely different books have been loaned in the two months. Values between 0 and 1 represent a partial reallocation of collective attention.
We distinguish between two main types of drift: local drift, which quantifies the JSD between a given month and the preceding month, and global drift, which measures drifts between a reference month and all other months (Fig. 1A). The two drift types reveal distinct patterns in how library book popularity evolves. While local drifts capture immediate shifts in book consumption, global drifts capture the longer-term changes to the loaning behavior.
Over the 4-year-span of our dataset, we observe variation in the number of library loans (Fig. 1B). In particular, two COVID-19-related lockdowns restricted library access in Denmark, resulting in a dramatic drop in how many physical books were loaned as well as an increase in digital book loans. Additionally, citizens tend to loan more books in the Summer holidays and less when libraries are typically closed (on Sundays and in the Winter holidays). To minimize noise and the effects of COVID-19, our analysis focuses on the period after the last lockdown (from May 2021 to December 2023). Individual book trajectories reveal considerable heterogeneity in how popularity evolves over time (Fig. 1C). While some books experience sharp peaks followed by steady decline, others maintain more stable popularity levels or exhibit complex patterns, including delayed and recurring peaks in popularity.
Collective consumer attention is highly unequally distributed among the 660,000 books in our dataset (Fig. 1D). An unequal popularity distribution is not unique to our dataset but is a commonly observed pattern across cultural markets [12, 13, 14, 15, 16]. In our data, most books are loaned less than 10 times, while only around 1,000 books attract more than 10,000 loans. Looking across book genres, loans appear to be most unequally distributed among adult fiction, and most equally distributed for nonfiction books (see inset in Fig. 1D). The shape of the popularity distributions is constant across months (see SI Fig. LABEL:book_popularity_distributions_months). This means that drifts in collective attention between two months capture changes in which books hold which positions in the rank-frequency distribution rather than simply changes in the shape of the distribution, such as increased or decreased diversity in book loans.
Loaner subgroups are not equally represented in this dataset (Fig. 1E). For instance, there is a clear gap between the sexes, with female loaners lending more books. Additionally, book types are not evenly distributed among loaner subgroups, especially among different age segments. Unsurprisingly, young loaners (0-17 years) almost exclusively loan children’s books. Loaners aged 30 to 45 take out many children’s books and relatively few adult fiction books, whereas 65+ loaners borrow very few children’s books and many adult fiction books. In SI Fig. LABEL:book_popularity_distributions_loaner_groups, we show the distribution of loans across different book types for each subgroup. Further, SI Fig. LABEL:book_popularity_distributions_loaner_groups_months demonstrates that these distributions are stable over time.
1.1 Patterns of collective drifts
We start our analysis at the most general level by quantifying drifts in popularity distributions across all books and loaners. Later, we will segment our analysis and investigate the contribution of different book categories and population demographics to drift. Fig. 2A shows the temporal dynamics of local and global drifts. Local drifts are approximately constant between successive months, punctuated only by small recurring peaks from November to January (red, solid line Fig. 2A). These deviations stem from season-specific books (predominantly Christmas-themed books) becoming popular in November and December. Global drifts, however, grow continually from the reference month of May 2021 (blue, solid line Fig. 2A), and show similar season-specific bumps. Taken together, this suggests two primary mechanisms of drift: a gradual, persistent shift in loaner preferences and a seasonal component causing momentary increases in collective attention.
Figure 2: Local and global drifts in cultural attention. A, Local (red) and global (blue) drifts across all book categories and loaner demographics. We estimate JSD in bits to ensure that drifts are bounded between 0 and 1. B-I, Local and global drifts split up according to population demographics and book category. To minimize the effects of random fluctuations and noise, we focus only on the top 10,000 most popular books. B-E, shows drifts for the top 10,000 most popular books, while panels F-I show drifts for the top 10,000 most popular adult books. Shaded areas around lines (barely visible) show $2\sigma$ bootstrap errors.
To explore whether patterns are similar across different book categories, we compute local and global drifts for three broad categories of books, which are available in our dataset: adult fiction, adult nonfiction, and children’s books. For all book categories, local drifts are stable, meaning that collective attention has a constant rate of drift, while global drifts grow over time (dashed and dotted lines Fig. 2A). However, drift patterns differ in notable ways. Specifically, adult nonfiction books have a larger rate of local drift compared to adult fiction and children’s books. One explanation for this is that a large proportion of nonfiction books are rarely loaned, as seen in Fig 1C. When we exclude infrequently loaned books, the local drift rate for nonfiction books becomes less extreme, but it is still higher than drifts in fiction and children’s books (see SI Fig. LABEL:drifts_top10k). For global drift, we observe different dynamics. Drift in children’s books grows at a considerably slower rate than in adult books. Lastly, for both local and global drifts, the seasonal peaks (around Christmas) are most pronounced for children’s books, less so for adult nonfiction books, and nearly absent for adult fiction books. In the SI, we show that these patterns are qualitatively robust: across different pairs of months (SI Fig. LABEL:drifts_heatmap_all_books-LABEL:drifts_heatmap_children), using generalized measures of drift (see SI Sec 1.1 and SI Figs. LABEL:drifts_alternative_measures_all_books-LABEL:drifts_alternative_measures_children), and for different temporal aggregation levels (SI Figs. LABEL:drifts_week-LABEL:drifts_quarter_top10k). We further show that patterns are qualitatively robust, although drift magnitudes are larger, across COVID lockdowns when we include data from January 2020 (SI Figs. LABEL:drifts_monthly_from-2020-LABEL:drifts_across-covid_comparison).
1.2 Drifts split up according to population subgroups
So far, we have investigated drifts across the entire Danish population. Here we investigate how dynamics vary across demographic subgroups. Prior work has shown that sociodemographic groups, defined by sex, age, education level, and area of residence to exhibit differences in the diversity of cultural consumption [38, 39]. Yet, to our knowledge, no study has quantified how the collective consumption of these segments evolves over time. To compare different subgroups of the loaner population, we account for the fact that drifts computed from raw frequencies capture both changes to the underlying collective attention and random fluctuations in the popularity of less popular books. To eliminate the effects of random fluctuations, we compute drifts exclusively for the top 10,000 most popular books within each sociodemographic subgroup and correct for any residual bias in drift estimation using bootstrap methods [40] (see Methods for details). For the comparison of loaners in different age bins, we exclude loaners below the age of 30 as loaners from 0 to 17 almost exclusively lend children’s books, and loaners between 18 and 29 have very few loans.
We find substantial differences in drifts between loaners of different ages (Fig. 2B). Both local and global drifts are considerably lower for 30-45-year-old loaners than for older ones, and the seasonal component is stronger. For sex (Fig. 2C), we find females have a larger global drift and stronger recurring seasonal peaks, while differences in local drift are minuscule. Additionally, it is worth noting that the frequency of seasonal peaks differs between males and females. While females show seasonal peaks centered around Christmas, males, in addition, also have summer peaks. For other sociodemographic subdivisions (education level and area of residence), differences in local drifts are not pronounced, and we observe only small to moderate differences between global drifts over time (Figs. 2D-E). Interestingly, seasonally recurring peaks are more pronounced for 30-45-year-olds and females, this is partially due to these groups loaning more children’s books (see Fig. 1E).
To investigate the role of children’s books in drift for individual demographic groups, we compare the results from Fig. 2B-E with drifts specifically calculated for the 10,000 most popular adult books (i.e., disregarding loans of children’s books). Here, we observe that the differences in global drifts between age groups are smaller when we only consider drifts in adult books (Fig. 2F). In addition, disregarding children’s books makes the differences in global drift more prominent between sex-, education-, and residence area-groups (Fig. 2G-I). Here we observe that drift is higher for females, for more educated individuals (Bachelor or higher), and for individuals living in large cities. These differences are robust when books are further split up into adult fiction and adult nonfiction (SI Fig. LABEL:drifts_subgroups_adult_fic_nonfic). When it comes to seasonal peaks, we find that removing children’s books eliminates the regularly occurring peaks for 30-45-year-olds (Fig. 2F), and both sexes (Fig. 2G). This is the case for both global and local drifts, and aligns well with younger demographics being more likely to have children.
1.3 Contributions from specific books
Until now, we have considered books collectively; here, we untangle the importance of individual items on drift and study what role individual books play in collective attention drift. We do this as much of the focus in both media and previous studies [1, 2, 3, 4, 8, 9, 25, 6] is on big hits that occupy our collective attention at any given moment; however, it is unclear how important individual items are for explaining overall drifts in collective attention. To investigate this, we measure how much each book contributes to (nation-scale) local and global drifts. For each month, we divide all books into five broad groups defined by the size of their contributions (see Methods for details). I.e., the more popular a book is, and the more its relative popularity fluctuates between months, the higher its contribution (see Methods). Groups are non-overlapping and defined as: i) the top 100 most contributing books, ii) the top 101 to 1000 most contributing books, iii) top 1001 to 10.000, iv) top 10.001 to 50.000, and v) books beyond top 50.000. Using these groups, we explore how many books we would need to account for a given proportion of the drift over time and how consistent the contributions of individual books are.
We start by examining how much each group of books contributes to the drift in collective attention over time (Fig. 3). Surprisingly, we find that the 100 books with the most significant shifts in their popularity consistently account for only around 10% of both local and global drifts (Fig. 3A-B). As such, even if it were possible to identify these books and predict changes in their popularity, 90% of the total shift in collective attention would remain unexplained, unless less influential books were also included. For local drift, the top 10,000 books that contribute most to drift account for less than 50% of the total change in behavior. This shows the importance of considering the full ‘cultural system’ and not just individual product trajectories when analyzing patterns of consumer attention. When it comes to seasonal variation in local drift (Fig. 3A) we observe increases in the relative contributions of the top 101-1,000 most increases, but the top 100 books contribute less to this seasonal pattern. For global drift, we find that the relative contributions of the top 1,000 books decline over time (Fig. 3B). In other words, the 1,000 books with the largest changes in popularity are less important in explaining collective attention drifts for months, the further we move from the reference month. Instead, shifts in the popularity of especially the top 1,001-10K most contributing books account for a larger proportion of global drifts over time. Nonetheless, even when accounting for the top 10.000 most influential books, we explain less than 70% of global drift over 2.5 years. This illustrates that books in the long tail of the popularity distribution, in a collective fashion, are impactful in driving changes in collective attention. In other words, when analyzing the cultural popularity of items, it is vital to consider books across all contribution groups.
Figure 3: Contributions of individual books to drift dynamics. A-B, Distributions of relative contributions of individual books to local drift (A) and global drift (B) over time. C, Probability that a book with a given level of contribution to the local drift in a month will have a certain level of contribution to the local drift in the next month. Probabilities are averaged across all pairs of months. D, Monthly popularity for the 1,000 books with the highest contributions to the global drift in October 2023, sorted by the month where their popularity peaks.
Although local drift is quite consistent across months, it is unclear to what extent books transition between the different contribution groups (top 100, top 101-1000, etc.) over time. To explore this, we measure the likelihood that a book belonging to a particular contribution group in a given month would remain in the same group or transition to one of the other groups in the following month. We average these transition probabilities over all pairs of consecutive months, and find that only 15% of the 100 books that contribute most to local drift in one month are also part of the 100 highest contributors in the next month (Fig. 3C). Taken together, individual books show substantial variation in their contributions to collective attention drift between consecutive months
To understand which books drive long-term shifts in collective attention, we examine the 1,000 books contributing most to global drift in October 2023 – the last month unaffected by the observed seasonal patterns – and order them by when they reach peak popularity (Fig. 3D). Two patterns emerge. First, rather than distributing evenly across our observation period, these books cluster at the boundaries of our observation window: many peak shortly after the reference month (May 2021), while others peak approaching October 2023. This clustering reveals that global drift arises from two distinct sources: books whose popularity has faded since early in our study period and those gaining recent prominence. Second, the popularity of these high-contributing books is remarkably short-lived. These two patterns – clustering at the edges of the observation period and volatile popularity trajectories – distinguish books driving the global drift from the most popular books and books with highest peaks in popularity (SI Figs. LABEL:popular_overall_counts-LABEL:peak_monthly_popular_counts), revealing that standard measures of book popularity do not fully capture which titles drive long-term collective attention shifts.
In summary, we find individual books to have limited power in explaining drifts in collective attention. Put differently, it is not possible to isolate a few books that account for drifts in a given month, even for drifts in the long term. As such, even if we had perfect predictions of the popularity trajectories of the books with the largest changes in popularity, they would only account for a small percentage of the total drift in collective attention.
1.4 Predicting future drifts from past behavior
So far, we have shown that the collective attention of library loaners is constantly changing. This has a direct impact on the predictability of the system. Further, it implies that the performance of any machine learning model fitted on the cultural consumption at a given point in time, e.g., to recommend books, will deteriorate over time if not updated and retrained. However, since drifts in collective attention exhibit stable patterns, it may be possible to predict attention drift and get an estimate on how fast model performance will decay and how often models should be retrained.
Here, we test the predictability of drifts in collective attention by predicting drifts in 2023 using information about drifts from 2022. As previously observed, drift in collective attention for library loaners is characterized by two components: a steady turnover and seasonal effects. If both of these components are stable across years, it is possible to predict the magnitude of drift in 2023 from the drift observed in 2022. To test this, we make predictions for both local and global drift. When predicting global drift, we use January 2022 and 2023 as baselines, respectively, for quantifying drift in 2022 and 2023. To make predictions, we use the following minimal model: for a given month in 2023, we use the observed (local and global) magnitude of drift for the same period in 2022 as the prediction. For example, if attention has drifted 0.2 bits between two months in 2022, we predict attention to drift with an equal amount between the two months in 2023.
Fig. 4 shows the observed and predicted drifts. Overall, this simple model performs exceptionally well, with minor deviations between the predicted and observed values. Performance remains strong when making predictions for all books (Fig. 4A), or when books have been segmented into different categories (Fig. 4B-D). Further, in SI Fig. LABEL:predictions_subgroups, we report predictions for all combinations of book and loaner subsets and find that the simple model also performs well for all these subsets. This showcases that although it is difficult to predict which items attention will drift towards, predicting the magnitude of drift is possible. This has important implications for how often algorithmic systems that recommend cultural items to users should be updated.
Figure 4: Predictions of future drifts from past drifts. A-D, shows the local and global drifts across all books (A) and for the three main subsets of books B-D, for 2022 and 2023. For each year, global drifts are computed with January as the baseline month.
2 Discussion
For decades, studies of cultural markets have emphasized the difficulty of predicting the ever-changing attention of cultural consumers [41, 42, 17, 18, 43]. We show that changes in the collective attention of Danish book loaners are consistently high and cannot be attributed to only a few books. Further, our study shows that the collective redistribution of attention across books in Danish libraries has identifiable patterns that can be predicted with a high degree of accuracy, even for specific groups of products and audiences. While we examine only broad product and audience groups, our findings suggest that predictable collective attention dynamics could exist for more fine-grained groupings, such as specific sub-genres.
Our findings raise important questions about the mechanisms underlying collective cultural change, specifically processes of individual choice and social influence. Relationships between individual choice, social influence, and cultural change have previously mainly been investigated in simulation studies [44, 45, 46, 47] and experimental work [17, 18, 48, 49, 50, 51]. Detailed registry data on individual loaners’ choices, such as the one we used in this study, can open new opportunities to study relationships between behavior at the individual-, group-, and macro-levels. A well-developed understanding of these mechanisms could be used to explain differences in how collective cultural attention evolves for different types of products and groups of consumers.
Our dataset contains comprehensive records of all Danish public library loans; however, our findings are not necessarily guaranteed to generalize across all book consumption. Book consumption is most likely different when there is an associated economic cost. For example, since borrowing books at the library is free, users may explore titles more widely than when they have to buy new books in a bookstore. Similarly, libraries feature a wide selection of both new and older titles, whereas bookstores prioritize newer releases and popular titles. These factors likely influence the drift of collective attention, but investigating their impact is beyond the scope of this work. Additionally, although we aggregate various media formats (physical books, ebooks, and audiobooks) to maximize coverage, consumption patterns may differ across these formats. Future work should look into this.
The extent to which our results can be generalized to other cultural markets is an empirical question; however, we expect similar forms of collective attention dynamics to apply widely. For instance, other cultural markets display variations of the drift components we observe for library loans: a stable turnover in consumer preferences and season-specific item consumption [52, 53]. Nonetheless, the importance of individual products for drift dynamics could differ significantly between markets. In markets with fewer products and more unequal distributions of popularity, a small number of products could account for a larger proportion of drift. In such markets, shocks by individual products could disrupt otherwise stable drift patterns, causing more unpredictable drift dynamics than what we observe for library loans. We hope that future studies will explore how characteristics of drifts in collective attention compare across cultural markets.
Recommender systems are another potential source of variation in consumer attention dynamics between markets. Recommender systems are pervasive in many current cultural markets, and their effects on drifts in collective consumer attention should be studied more closely. However, because these systems are so deeply embedded in most real cultural markets, isolating their effects is a challenge [54]. Even public libraries rely heavily on centralized recommendation systems, both on their websites for book search and reservations, and via librarians’ curation of books in each local library.
Our findings have major implications for adapting to changing consumer attention. With reliable expectations of drifts in consumer attention, individuals and organizations interested in predicting and modeling cultural consumption can make more informed decisions about when machine learning models require updating. As we show, drifts in collective attention can be faster for certain product types and consumer groups (Fig. 2). Model updating should take into account these group-specific drift differences, as failing to do so could result in problematic biases towards groups with faster drifts in collective attention. For example, drift is higher for females and for 65+ year demographics, meaning models need to be updated more frequently to accurately reflect their tastes. The consequences of drifts in collective attention will depend on the specific model. In particular, many recommender systems are based on collaborative filtering approaches that capture associations between users and items [55]. In these models, performance declines can result not only from drift in item popularity but also from drift in the user base and drift in the association between users and items. Our study focuses solely on item popularity, leaving open important questions about the predictability of drift in other data used to model cultural consumption.
We envision the methods developed in this study to be used to study changes in collective behavior in other domains outside of cultural markets. Like cultural consumption, behaviors that are socially driven are generally difficult to predict, including information cascades, social influence, and social interactions [56, 57, 58, 59]. For such many behaviors, it may be possible to make higher-order predictions of behavior that can be used to contextualize and adapt to uncertain predictions of individual phenomena like the popularity of specific topics or items.
3 Methods
Data Access & Anonymity. All data processing and analysis was conducted in-house at Statistics Denmark, which is the national statistical agency of the country, in accordance with their data protection protocols. This means that all individual-level data, both the library takeout dataset and the sociodemographic loaner information, are pseudonymized and kept strictly confidential. Due to the sensitive nature of the data, it is not possible to openly share the data, but it is possible to apply for access to the data through Statistics Denmark.
The Danish Nation-Scale Library Takeout Dataset. The library dataset used in this study is based on administrative registers of library takeouts from Danish public libraries, collected by Statistics Denmark. The dataset is continuously updated, and the version used in this study spans from January 1, 2020, to December 31, 2023, and includes all loans from public libraries, including Ereolen (the Danish national platform for digital books), but excluding university and school libraries.
Sociodemographic Loaner Information. All variables except sex are dynamic in our dataset. This means that loaners can belong to different subgroups at different points in time. Below, we describe how we derived the different loaner categories. Age is computed directly from each loaner’s birthdate at the time of each loan. Residence Area is based on standard definitions provided by Statistics Denmark, which divides municipalities into five types (https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper), which we collapse into two main types: Large city municipality, which is defined by a municipality with above 100,000 citizens or has a high job availability (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper for specifics), and Town or rural municiplaity, which covers all other municipalities. Level of Education is recorded as the loaner’s maximum attained education level at the time of each loan. The three categories are based on the DISCED-15 classification (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/disced15-udd), which groups education into 15 main categories.
Library Data Preprocessing.
Item Filtering. Around 10 million loans in the dataset refer to non-books (see SI Fig. LABEL:material_type_distribution for the distribution of loans across material types), which were removed from analysis.
Book Identification. Each item is associated with a unique identification number, known as a faust number. Because each edition of a book gets a separate faust number, one title can have several associated faust numbers. Since we are interested in the evolution in the popularity of titles, regardless of the medium (physical book, audiobook, etc.), we treat different media variants of the same book as a single entity. To derive a unique identifier independent of the edition and medium, the following heuristic algorithm was applied:
  1. 1.
    Remove special characters from book titles, and make all characters in the titles lowercase.
  2. 2.
    Sort all book titles and authors lexicographically.
  3. 3.
    For each book, compare it with the succeeding ten books in the sorted list, and find all pairs of books that satisfy the following criteria:
    1. (a)
      The edit distance between their titles and authors is at most 1.
    2. (b)
      Both book titles contain the same digit of maximum length 2, indicating a particular version of the book. Book titles with no digits and books with a digit of 1 in the title are treated as the same (e.g., Ternet Ninja and Ternet Ninja 1).
  4. 4.
    For all pairs that follow these criteria, group together any pair where one of the books overlaps.
This process reduced the unique number of items from approximately 784,000 to around 660,000.
Book Category Book categories (adult fiction, adult nonfiction, children’s) were provided by the Danish Bibliographic Centre and not further preprocessed.
Measuring Changes in Collective Consumer Attention.
Jensen-Shannon Divergence and Book-Level Contributions. In each month, we compute the popularity (number of loans) of each book $i$ relative to the total amount of loans in that month. The drift in collective attention between those months is the JSD between the distributions of popularity in two time months $P$ and $Q$, which can be expressed as a linear combination of changes to the popularity of each individual book:
$\text{JSD}(P,Q)$ $=\sum_{i}\Bigl(\overbrace{{p_{i}\log\frac{2p_{i}}{p_{i}+q_{i}}}+{q_{i}\log\frac{2q_{i}}{p_{i}+q_{i}}}}^{\text{Partial JSD}_{i}}\Bigr)$ (1)
where $p_{i}$ is the relative popularity of book $i$ in the first month, $q_{i}$ is the relative popularity of book $i$ in the second period. The partial JSD represents the contribution of each book to the overall JSD. If there is no change to a book’s relative popularity, it does not contribute to the JSD. For any measure of drift, the relative contribution of a group of books to that drift is simply calculated as the sum of partial JSDs for books in that group divided by the JSD.
Estimation of Jensen-Shanon Divergence. Empirical estimation of information-theoretic quantities from samples is known to suffer from bias, particularly when the number of categories (here: unique books) is larger than the total number of sampled counts (here: loans) [60]. To maintain a consistent relation between changes in overall preferences and changes in preferences for individual books, we use the maximum likelihood estimator when computing JSDs for all loaners.
To ensure robust comparisons between sociodemographic subgroups, we compute JSDs only for the top 10,000 books in each group. We use a bootstrap estimator, which has been shown to outperform other estimators in cases where the number of counts is higher than the number of categories [40]. Additionally, this estimator provides uncertainty estimates for each computed JSD. 500 bootstrap permutations were used to compute each bootstrap estimate.
4 Acknowledgements
We are grateful for the help with data collection by the Data Science Lab at Statistics Denmark and Ea Hoppe Blaabæk.